# Zero-shot Safety Prediction for Autonomous Robots with Foundation World Models

### v2

@@ Abstract
A world model creates a surrogate world to train a controller and predict safety violations by learning the internal dynamic model of systems. However, the existing world models rely solely on statistical learning of how observations change in response to actions, lacking precise quantification of how accurate the surrogate dynamics are, which poses a significant challenge in safety-critical systems. To address this challenge, we propose foundation world models that embed observations into meaningful and causally latent representations. This enables the surrogate dynamics to directly predict causal future states by leveraging a training-free large language model. In two common benchmarks, this novel model outperforms standard world models in the safety prediction task and has a performance comparable to supervised learning despite not using any data. We evaluate its performance with a more specialized and system-relevant metric by comparing estimated states instead of aggregating observation-wide error.
@@ I Introduction
The rise of foundation models provides an opportunity to create meaningful causal representations with a zero-shot segmentation of observed images. Not only the causal representations can simplify the prediction and also the whole world model can also be implemented with a training-free architecture using Large Language Models, eliminating the need to collect and label training data.
This paper proposes foundation world models that further reduce data requirements by using foundation models in two key elements of world models. First, to obtain causal latent representations, we use the Segment Anything Model (SAM) [7] to get the pixel positions of all objects in the observation. The important characteristics of latent states, in particular their safety (e.g., whether a collision has occurred or not), can be calculated based on these representations. Second, we predict the future position of these objects with a Large Language Model (LLM). In addition, we introduce a more focused metric for the accuracy evaluation of predicted state. This metric helps us evaluate the quality of surrogate dynamics in a system-specific way.
Our evaluation on two common simulated benchmarks demonstrates the value of foundation world models. Our foundation model not only demonstrates superior state prediction based on the newly proposed metric but also excels in safety predictions. We experiment with several baseline approaches, evaluation metrics, and different Large Language Models for latent prediction.
@@
    A training-free world model that combines foundation models with causally meaningful embedding and overcomes the distribution shift of the predicted observation, which occurs in standard world models.
@@ II-B Object-based Prediction Metric
For evaluating the image-level error between a prediction $\hat{\bm{y}}$ and a ground truth $\bm{y}$, we will implement the segmentation algorithm to split the image into several parts: $\Omega({\bm{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...,{\bm{\omega}}_{n}]$, $\Omega({\hat{\bm{y}}})=[\hat{\bm{\omega}}_{1},\hat{\bm{\omega}}_{2},...,\hat{\bm{\omega}}_{n}]$ get a vector of CDs: $[\text{CD}(\bm{\omega}_{1},\hat{\bm{\omega}}_{1}),\text{CD}(\bm{\omega}_{2},\hat{\bm{\omega}}_{2}),...,\text{CD}(\bm{\omega}_{n},\hat{\bm{\omega}}_{n})]$ and can precisely reflect how well each part is predicted instead of providing all errors of unrelated pixels.
@@ II-C Safety Prediction Problem
Predicate $\varphi$ is a binary function of state $\bm{x}_{t}$ that determines the safety of the system at time $t$: $\varphi(\bm{x}_{t})$. For example, it can indicate if the robot is dangerously close to an obstacle. The safety prediction problem [5] is defined as follows:
@@
A significant challenge in standard world models is that the safety is not straightforward to infer from the high-dimensional observation, so it necessitates an extra learning-based safety evaluator (often a Convolutional Neural Network, CNN) shown on the left of Fig. 2. This CNN safety evaluator (as well as the controller) can suffer from distribution shift due to distorted predictions $\hat{\bm{y}}$ of standard world models originating in the decoder. Specifically, the conditional distribution of safety is the same for training and test datasets $p_{train}(\varphi(\bm{x})\mid\bm{y})=p_{test}(\varphi(\bm{x})\mid\bm{y})$, but the distribution of observations is shifted: $p_{train}(\bm{y})\neq p_{test}(\bm{y})$, which influences the accuracy of the CNN evaluator.
@@
10:         $\bm{z}\leftarrow\bm{z}+\bm{z}_{i}^{j}$triangleright Store the centroids for calculating the error later
@@ III Foundation World Models
8:         SWAP($\bm{y}_{t+m+1}[px][py]$, $\bm{y}_{t+m+1}[px+{\delta}^{N}_{x}][py+{\delta}^{N}_{y}]$) $\triangleright$ swap the pixel values
@@
Method Input length Horizontal position error Vertical position error
range: [0,20] range: [0,20]
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@
TABLE I: State prediction performance for the lunar lander: the horizontal and vertical centroid distance (CD) errors. Gray rows indicate the use of additional data.
Method Input length Position (pixel range: [-48,48]) Angle (degree range [-180°,180°])
@@
Method Input length F1 score $\uparrow$ FPR $\downarrow$
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@ IV-A Experimental Setup
Figure 4: The observation of cart pole (upper) and lunar lander (lower). From left to right, generated by: the true observation model, a foundation world model, an existing world model in-distribution, and an existing world model out-of-distribution.
@@ IV-B Experimental Results
Method Input length SSIM $\uparrow$ MSE $\downarrow$
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@ V-A Foundation Models and Applications
With the rise of large AI models that benefit from the availability of huge computing resources by using a considerable amount of data and model parameters, foundation models achieve great performance on multiple cross-domain tasks. Many well-performing models like Generative Pre-trained Transformer (GPT) [14], Bidirectional Encoder Representations from Transformers (BERT) [15], Contrastive Language-Image Pre-Training (CLIP) [16] and Segment Anything model (SAM) [7] extensively serve as the foundation for downstream tasks. The core idea of Large Language Models is to capture statistical patterns of language by learning from large amounts of text data rather than truly understanding the meaning or logic of the content.
Large language models can also complete complex sequence tasks that are not limited to text but extend — without specialized training — to domains such as robotics [17]. In addition to applying to decision-making, language-guided abstraction [18] can transfer high-level task descriptions into task-relevant state abstractions by using a pre-trained language model. Abstract meaning representation also plays an important role in LLMs and can improve their performance [19]. Since human and robotic systems may have different understandings of natural languages, human and robot representation alignment is performed to close this gap [20].
As the embedded structure of LLM is a generative pre-trained transformer, it can be used for prediction. Some researchers have tested LLM’s forecasting ability with humans and it shows that the ability of LLM is approaching the human level [21]. Some new foundation models like Chronos [22] are specialized to execute non-dynamical prediction tasks like traffic, weather, and energy consumption. Language can also be the input of robotic systems [23] to be trained for end-to-end language-controlled systems like RT2 [24], which can perform basic semantic reasoning to finish its tasks. The trustworthiness of robotic techniques based on foundation models is still an area that has not been widely explored [25]. Besides applying the foundation model to control and forecast, some novel simulator also adopts LLM to generate various scenes for traffic simulations [26].
@@ V-B Safety Assurance of Autonomous Systems
Trajectory forecasting plays a key role within autonomous systems for a variety of purposes, including safety assurance. For sensors operating in high-dimensional spaces, integrating physics models is crucial for enhancing predictive accuracy. Works such as the Social ODE [27] and Deep Kinematic Models [28] exemplify this approach by merging deep learning techniques with physical models to address issues related to data unreliability. Another popular safety assurance method is reachability analysis. Given the challenge of finding all reachable states in non-linear dynamics and neural network controllers, specialized verification tools [29, 30] are designed to develop accurate overapproximations of these reachable sets. However, these tools cannot work on high-dimensional controllers such as image-input controllers. High-dimensional verification processes could be tailored, for instance, through the application of generative models [31], and by approximating high-dimensional systems with their lower-dimensional counterparts [32]. Finally, conformal prediction is also widely recognized for establishing boundaries around time series forecasts within a defined confidence level  [33, 34], which could be used to provide safety bounds for the trajectory with a specific confidence.

### v3

@@ Abstract
A world model creates a surrogate world to train a controller and predict safety violations by learning the internal dynamic model of systems. However, the existing world models rely solely on statistical learning of how observations change in response to actions, lacking precise quantification of how accurate the surrogate dynamics are, which poses a significant challenge in safety-critical systems. To address this challenge, we propose foundation world models that embed observations into meaningful and interpretable latent representations. This enables the surrogate dynamics to directly predict interpretable future states by leveraging a training-free large language model. In two common benchmarks, this novel model outperforms standard world models in the safety prediction task and has a performance comparable to supervised learning despite not using any data. We evaluate its performance with a more specialized and system-relevant metric by comparing estimated states instead of aggregating observation-wide error.
@@ I Introduction
The rise of foundation models provides an opportunity to create meaningful representations with a zero-shot segmentation of observed images. Not only the interpretable representations can simplify the prediction and also the whole world model can also be implemented with a training-free architecture using Large Language Models, eliminating the need to collect and label training data.
This paper proposes foundation world models that further reduce data requirements by using foundation models in two key elements of world models. First, to obtain interpretable latent representations, we use the Segment Anything Model (SAM) [7] to get the pixel positions of all objects in the observation. The important characteristics of latent states, in particular their safety (e.g., whether a collision has occurred or not), can be calculated based on these representations. Second, we predict the future position of these objects with a Large Language Model (LLM). In addition, we introduce a more focused metric for the accuracy evaluation of predicted state. This metric helps us evaluate the quality of surrogate dynamics in a system-specific way.
Our evaluation on two common simulated benchmarks demonstrates the value of foundation world models. Our foundation model not only demonstrates superior state prediction based on the newly proposed metric — but also excels in safety predictions. We experiment with several baseline approaches, evaluation metrics, and different Large Language Models for latent prediction.
@@
    A training-free world model that combines foundation models with interpretable embedding and overcomes the distribution shift of the predicted observation, which occurs in standard world models.
@@ II-B Object-based Prediction Metric
For evaluating the image-level error between a prediction $\hat{\bm{y}}$ and a ground truth $\bm{y}$, we implement the segmentation algorithm to split the image into several parts: $\Omega({\bm{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...,{\bm{\omega}}_{n}]$, $\Omega({\hat{\bm{y}}})=[\hat{\bm{\omega}}_{1},\hat{\bm{\omega}}_{2},...,\hat{\bm{\omega}}_{n}]$. Then, we get a vector of CDs $[\text{CD}(\bm{\omega}_{1},\hat{\bm{\omega}}_{1}),\text{CD}(\bm{\omega}_{2},\hat{\bm{\omega}}_{2}),...,\text{CD}(\bm{\omega}_{n},\hat{\bm{\omega}}_{n})]$ that precisely reflects how well each object is predicted, rather than aggregating the errors of unrelated pixels. We leave the cases of missing and ghost objects for future work.
@@ II-C Safety Prediction Problem
The safety predicate $\varphi$ is a binary function of state $\bm{x}_{t}$ that determines the safety of the system at time $t$: $\varphi(\bm{x}_{t})$. For example, it can indicate if the robot is dangerously close to an obstacle. The safety prediction problem [5] is defined as follows:
@@
A significant challenge in standard world models is that the safety is not straightforward to infer from high-dimensional observations, so it necessitates an extra learning-based safety evaluator (often a Convolutional Neural Network, CNN) shown on the left of Fig. 2. This CNN safety evaluator (as well as the controller) can suffer from distribution shift due to distorted predictions $\hat{\bm{y}}$ of standard world models originating in the decoder. Specifically, the conditional distribution of safety is the same for training and test datasets $p_{train}(\varphi(\bm{x})\mid\bm{y})=p_{test}(\varphi(\bm{x})\mid\bm{y})$, but the distribution of observations is shifted: $p_{train}(\bm{y})\neq p_{test}(\bm{y})$. This kind of distribution shift is covariate shift [8], which influences the accuracy of the CNN evaluator.
@@
10:         $\bm{z}\leftarrow\bm{z}+\bm{z}_{i}^{j}$
@@ III Foundation World Models
8:         SWAP($\bm{y}_{t+m+1}[px][py]$, $\bm{y}_{t+m+1}[px+{\delta}^{N}_{x}][py+{\delta}^{N}_{y}]$) $\triangleright$ Move the object into the predicted position
@@
Horizontal position CD error Vertical position CD error
Method Input length k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@
TABLE I: State prediction performance for the lunar lander: the horizontal and vertical centroid distance (CD) errors. Gray rows indicate the use of additional data. The range of position is [0, 20].
Method Input length Position CD error (pixel range: [-48, 48]) Angle (degree range [-180°, 180°])
@@
F1 score $\uparrow$ FPR $\downarrow$
Method Input length k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@ IV-A Experimental Setup
Figure 4: The observation of cart pole (upper) and lunar lander (lower). From left to right, generated by: the true observation model, a foundation world model, an existing world model without distribution shift, and an existing world model with distribution shift.
@@ IV-B Experimental Results
SSIM $\uparrow$ MSE $\downarrow$
Method Input length k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
@@ V-A Foundation Models and Applications
With the rise of large AI models that benefit from the availability of huge computing resources by using a considerable amount of data and model parameters, foundation models achieve high performance on multiple cross-domain tasks. Many well-performing models like Generative Pre-trained Transformer (GPT) [14], Bidirectional Encoder Representations from Transformers (BERT) [15], Contrastive Language-Image Pre-Training (CLIP) [16] and Segment Anything model (SAM) [7] extensively serve as the foundation for downstream tasks. We extend the core idea of LLMs from text prediction to complex sequence tasks such as safety prediction, by exploiting the statistical patterns of language learned from large text data.
Furthermore, one significant advantage of LLMs is their extensive pre-training, making them easily adapted to various domains such as robotics decision-making tasks [17]. Another language-guided abstraction [18] can transfer high-level task descriptions into task-relevant state abstractions by using a pre-trained language model. Abstract meaning representation also plays an important role in LLMs and can improve their performance [19]. We believe that LLNs can achieve better performance on complex task sequences because the alignment between human and robot representations is closer than that of human and robot understanding [20].
As the internal structure of LLM is a generative pre-trained transformer, it can be used for any prediction probelms. Some researchers have tested LLM’s forecasting ability with humans and it shows that the ability of LLM is approaching the human level [21]. Some new foundation models like Chronos [22] are specialized in non-dynamical prediction tasks like traffic, weather, and energy consumption. Language can also be the input of robotic systems [23] to be trained for end-to-end language-controlled systems like RT2 [24], which can perform basic semantic reasoning to finish its tasks. The trustworthiness of robotic techniques based on foundation models is still an area that has not been widely explored [25]. Besides applying the foundation model to control and forecast, novel simulators also adopt LLMs to generate various scenes for traffic simulations [26].
@@ V-B Safety Assurance of Autonomous Systems
Trajectory forecasting plays a key role within autonomous systems for a variety of purposes, including safety assurance. One tricky problem with safety assurance based on trajectory prediction is the gap between high-dimensional input and low-dimensional states. For sensors operating in high-dimensional spaces, integrating physics models is crucial for enhancing predictive accuracy. Works such as the Social ODE [27] and Deep Kinematic Models [28] exemplify this approach by merging deep learning techniques with physical models to address issues related to data unreliability. Another popular safety assurance method is reachability analysis. Given the challenge of finding all reachable states in non-linear dynamics and neural network controllers, specialized verification tools [29, 30] are designed to develop accurate overapproximations of these reachable sets. However, these tools cannot work on high-dimensional controllers such as image-input controllers. One tricky problem with safety assurance based on trajectory prediction is the gap between high-dimensional input and low-dimensional states. High-dimensional verification processes could be tailored, for instance, through the application of generative models [31], and by approximating high-dimensional systems with their lower-dimensional counterparts [32]. Unlike the cases above, we apply SAM to bridge the gap between high-dimensional and low-dimensional space, which requires no additional training like LLNs and is easy to implement.
Finally, conformal prediction is also widely recognized for establishing boundaries around time series forecasts within a defined confidence level  [33, 34], which could be used to provide safety bounds for the trajectory with a specific confidence.